OpenStudy (anonymous):

Solve the following equation: (1/x)+2/(x-1)=3

OpenStudy (anonymous):

cross multiply the left side of the equation.

OpenStudy (anonymous):

\[x=\pm \sqrt{(6x-1)/3}\]

OpenStudy (anonymous):

that will give you (x-1+2x)/(x^-x)

OpenStudy (anonymous):

so that is incorrect

OpenStudy (anonymous):

what I am saying is the first step to solving the equation.. I did not state the answer...

OpenStudy (anonymous):

oh, is the answer correct

OpenStudy (anonymous):

that is what you get from cross multiplication of the left side of your equation....

OpenStudy (anonymous):

kumar hasnt been wrong yet.. i would bet he is right

OpenStudy (anonymous):

i was just trying to teach you the problem..

OpenStudy (anonymous):

x(x-1)1/x+x(x-1)2/x-1=x(x-1)3 (x-1)+2x=3x(x-1) (x-1)+2x=3x^2-3x 3x-1=3x^2-3x

OpenStudy (anonymous):

Correct, so far?

OpenStudy (anonymous):

3x^2-3x-3x+1=0

OpenStudy (anonymous):

3x^2-6x+1=0

OpenStudy (anonymous):

how am I doing?

OpenStudy (anonymous):

close..

OpenStudy (anonymous):

step back to the beginning..

OpenStudy (anonymous):

when you cross multiply.. leave the right side of the equation alone..

OpenStudy (anonymous):

so just on the left side.. ( (x-1)(1) + 2(x) ) / (x * (x-1))

OpenStudy (anonymous):

and that simplifies to (3x-1) / (x^2-x)

OpenStudy (anonymous):

again... just on the left side

OpenStudy (anonymous):

you with me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. now multiply both sides to get rid of the denominator on the left...

OpenStudy (anonymous):

so you move the (x^-x) over

OpenStudy (anonymous):

and get 3x-1 = 3x^ - 3x

OpenStudy (anonymous):

so 3x-1=3x^2-3x

OpenStudy (anonymous):

xactly.

OpenStudy (anonymous):

thats what i had

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

3x^2-6x+1=0

OpenStudy (anonymous):

quadratic equation...

OpenStudy (anonymous):

i'm bad at factoring

OpenStudy (anonymous):

i got +-sqrt(5)

OpenStudy (anonymous):

shouldnt be any more factoring

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

should be able to plug into the quadratic equation......

OpenStudy (anonymous):

(-b+-sqrt(b^2-4ac)) / 2a

OpenStudy (anonymous):

what is the step after 3x^2-6x+1=0

OpenStudy (anonymous):

this...? (-b+-sqrt(b^2-4ac)) / 2a